Mission Operations Days: 2001/118 to 2001/124
April 28, 2001 00:00:00 GMT to May 04, 2001 00:00:00 GMT
BAND 5 GAIN CHANGES were made at 113/13:17:53 – 13:28:08.
The solar diffuser sector size was changed from 50 frames to 5 frames, and the SRCA sector size was changed from 10 frames to 55 frames, from 117/12:11:33 to 117/17:15:13. The SCIABNORM flag was set from 117/12:10:39 to 117/17:22:27.
488,682 CADU’s of MODIS Science Data Lost on OPS DOY 2001/117. See Non-MODIS Significant Events below for details.
Terra Spacecraft and MODIS Instrument Status:
|Terra (AM-1) is in Normal Ops Mode|
MODIS is in Science Mode - B Side
|Blackbody||A Off; B On (290 K)||Nominal|
|Calibration Electronics||A Off; B On||Nominal|
|Control Processor||A Off; B On||Nominal|
|Nadir Door||Unlatched, Open||Nominal|
|Space View Door||Unlatched, Open||Nominal|
|Solar Diffuser Door||Unlatched, Closed||Nominal|
|FDDI Formatter||A Off; B On||Nominal|
|FDDI Port||A On; B Off||Nominal|
|FIFO Memory||#1 & #2 Off; #3 & #4 On||Nominal|
|Format Processor||A Off; B On||Nominal|
|Power Supplies||#1 Off; #2 On||Nominal|
|PV VIS FPA||A Off; B On||Nominal|
|PV NIR FPA||A Off; B On||Nominal|
|PV SMIR FPA||A Off; B On||Nominal|
|PV LWIR FPA||A Off; B On||Nominal|
|PC LWIR FPA||A Off; B On||Nominal|
|RC Outgas Heaters||All Off (CS, IS, OS)||Nominal|
|RC LWIR FPA Heater||On (83 K)||Nominal|
|RC SMIR FPA Heater||Off||Nominal|
|Scan Assembly||A Off; B On||Nominal|
|SDSM||A Off; B Off||Nominal|
|SRCA||A Off; B Off||Nominal|
|PS1 Survival Heater||Enabled||Nominal|
|PS2 Survival Heater||Enabled||Nominal|
|Timing Generator||A Off; B On||Nominal|
|Flight Software||Rev BD + 2 set of patches||Nominal|
|Inhibit Ids Set||None||Nominal|
This Week’s Completed MODIS Activities:
Saturday, April 28, 2001
Sunday, April 29, 2001
Monday, April 30, 2001
Tuesday, May 01, 2001
Wednesday, May 02, 2001
Thursday, May 03, 2001
2001/123 16:02:31-16:13:59 ATC OA15: SD/SDSM OPEN
2001/123 17:41:23-17:52:51 ATC OA16: SD/SDSM SCREENED
Friday, May 04, 2001
This Week’s Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Saturday, May 05, 2001
Sunday, May 06, 2001
Monday, May 07, 2001
Tuesday, May 08, 2001
None (Drag Make Up Maneuver, see below)
Wednesday, May 09, 2001
Thursday, May 10, 2001
Friday, May 11, 2001
There will be no MODIS roll maneuver in May because the roll angle would have been too great (more than 20-degree roll required). The next attempt for a roll maneuver will be June.
The next Drag Make Up Maneuver is scheduled for Wednesday, May 8, 2001 (OPS DOY 2001/128). Time of the maneuver will be 20:20:30 with an18 second duration. The expected loss and re-acquisition of pointing accuracy period is 18:50:30-21:50:30.
General Instrument Comments:
MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).
MODIS Telemetry Trends:
Telemetry is trending nominally.
Non-MODIS Significant Events:
488,682 CADU’s of MODIS Science Data Lost on OPS DOY 2001/117. 2 scheduled TDRS contacts were deleted because of space shuttle priorities (01:17 and 01:47). An emergency TDRS contact at 01:45 was unsuccessful (failed lock on both I and Q channels). An attempt to schedule an emergency contact with AGS was unsuccessful (the S/C was not properly configured, and AGS could not guarantee the pass because of a software upgrade). At 117-02:49:48 the MODIS SSR Buffer filled and was disabled. The Buffer was re-enabled at 117-03:05:16 after completion of a playback.
Near 125/00:00 MISR was transitioned (via its internal commanding) to its SAFE mode during one of its activities. More information will be provided once the anomaly is further diagnosed.
Limited Life Item Status:
SRCA 10W Lamp #1: 185.5 of 500 hours
SRCA 10W Lamp #2: 142.4 of 500 hours
SRCA 10W Lamp #3: 151.6 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours
SRCA 1W Lamp #1: 557.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours
Solar Diffuser Door: 1603 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements